# SpectralEarth: Training Hyperspectral Foundation Models at Scale

### v1

@@ Abstract
Foundation models have triggered a paradigm shift in computer vision and are increasingly being adopted in remote sensing, particularly for multispectral imagery. Yet, their potential in hyperspectral imaging (HSI) remains untapped due to the absence of comprehensive and globally representative hyperspectral datasets. To close this gap, we introduce SpectralEarth, a large-scale multi-temporal dataset designed to pretrain hyperspectral foundation models leveraging data from the Environmental Mapping and Analysis Program (EnMAP). SpectralEarth comprises 538,974 image patches covering 415,153 unique locations from more than 11,636 globally distributed EnMAP scenes spanning two years of archive. Additionally, 17.5% of these locations include multiple timestamps, enabling multi-temporal HSI analysis. Utilizing state-of-the-art self-supervised learning (SSL) algorithms, we pretrain a series of foundation models on SpectralEarth. We integrate a spectral adapter into classical vision backbones to accommodate the unique characteristics of HSI. In tandem, we construct four downstream datasets for land-cover and crop-type mapping, providing benchmarks for model evaluation. Experimental results support the versatility of our models, showcasing their generalizability across different tasks and sensors. We also highlight computational efficiency during model fine-tuning. The dataset, models, and source code will be made publicly available.
1 Introduction
@@
Hyperspectral imaging (HSI) from space provides valuable information about the material composition of the Earth’s surface and atmosphere. With hundreds of narrow bands, each a few nanometers in width, hyperspectral images record detailed electromagnetic information across a wide range of wavelengths, from long-wave ultraviolet to the shortwave infrared [39]. This rich spectral information provides opportunities for numerous environmental applications such as soil and mineral mapping, pollution tracking, agricultural assessment, and forest monitoring [13, 62, 35, 45, 61].
@@
Figure 1: Visualization of scale for various hyperspectral datasets published within the past three decades. The plot serves to illustrate the volume of the SpectralEarth dataset (area of circles).
@@
In recent years, the availability of hyperspectral data has been significantly expanded by the launch of new satellite missions. Notable examples include: Germany’s Environmental Mapping and Analysis Program (EnMAP) [37] with precursor Earth Sensing Imaging Spectrometer mission (DESIS) [38] mounted to the International Space Station (ISS), Italy’s Hyperspectral Precursor and Application Mission (PRISMA) [52], and the European Space Agency’s upcoming Copernicus Hyperspectral Imaging Mission for the Environment (CHIME) [53]. These developments open new avenues to employ deep learning and Self-Supervised Learning (SSL) for large-scale HSI analysis.
@@
Figure 2: Mosaic of pseudo-RGB representatives from the SpectralEarth dataset to showcase a diverse range of landscapes including urban areas, agricultural land, deserts, forests, and water bodies. Each image is presented with a false-color composite using the first three principal components, highlighting the spectral diversity captured across environments.
@@
Foundation models pre-trained with SSL demonstrated remarkable generalization capabilities with minimal fine-tuning in computer vision [51]. Subsequently, there has been a surge in developing geospatial foundation models, particularly for multispectral and high-resolution RGB imagery [64]. This line of research benefits from the abundance of publicly available satellite data from missions like Copernicus Sentinel-2 which provides petabytes of archive data for model training. In contrast, progress in foundation models for the hyperspectral domain has been hindered by the lack of large-scale HSI datasets for pre-training. Historically, access to hyperspectral imagery has been costly, with most benchmarks restricted to small-scale aerial imagery datasets [3]. While recent efforts [19, 55, 63] have begun to address this limitation, they still fall short in terms of data volume and geographic diversity to effectively train general hyperspectral foundation models.
@@
To bridge this gap, we introduce SpectralEarth, a large-scale dataset derived from the EnMAP satellite mission, featuring over 538,974 hyperspectral image patches from 415,153 unique locations with global spatial distribution and cloud coverage below 10%. SpectralEarth, as illustrated in Figure 1, represents an important leap in scale, being significantly larger than existing HSI datasets, and spans a wide variety of landscapes to reflect the full diversity of spectral signatures. Compiled from 11,636 EnMAP scenes, the dataset volume exceeds 3TB in hyperspectral images. Notably, about 17% of its geospatial locations include timeseries data to enable multi-temporal HSI analysis. To effectively harness the rich information in hyperspectral data, we modify conventional vision backbones such as ResNet and ViT with a spectral adapter module. It enables such backbones to capture the unique spectral characteristics of HSI. Using three popular SSL algorithms, MoCo-V2 [26, 15], DINO [12] and MAE [27], we train these backbones on SpectralEarth, to provide a plurality of pretrained models for hyperspectral image analysis. To benchmark the efficacy of our models, we introduce four downstream datasets geospatially co-registered with EnMAP and DESIS imagery for land cover and crop type mapping. Our experiments demonstrate the potential of the foundation model paradigm to efficiently generalize across various hyperspectral imaging contexts, including adapting to data from alternative sensors with different spectral characteristics. Our contributions summarized:
@@
  • Assembly of SpectralEarth, a large-scale dataset comprising around 3.3TB of nearly cloud-free EnMAP hyperspectral images. SpectralEarth features a global geospatial distribution encompassing over 538,974 image patches from 415,153 unique geo-locations, 73,307 of which include multiple timestamps.
@@
  • Construction of four downstream datasets designed to benchmark classification and semantic segmentation tasks for HSI.
@@
  • Training of various foundation models on SpectralEarth to test and demonstrate their generalizability to multiple hyperspectral sensors. In addition, we demonstrate the models’ computational benefit through faster convergence in fine-tuning.
2 Related Work
@@ 2.1 Hyperspectral Datasets
Indian Pines 1 145x145 AVIRIS 224 20m $\times$
KSC 1 512x614 AVIRIS 224 18m $\times$
Salinas 1 512x217 AVIRIS 224 3.7m $\times$
Pavia Uni. 1 610x340 ROSIS 103 1.3m $\times$
Pavia Center 1 1096x1096 ROSIS 102 1.3m $\times$
Botswana 1 1476x256 EO-1 242 30m $\times$
Cooke City 1 280x800 HyMap 126 3m $\times$
Houston Uni. 1 349x1905 ITRES-CASI 144 2.5m $\times$
SDFC 1 1201x201 HAHS 63 0.5m $\times$
Chikusei 1 2517x2335 Hyperspec-VNIR-C 128 2.5m $\times$
WHU-Hi 3 1217x303, 940x475, 550x400 Nano-Hyperspec 270-274 0.043-0.463m $\times$
Hyperview $\sim$3k avg. $\sim$60x60 HySpex 150 2m $\times$
@@
MSST $\sim$20k 64x64 EnMAP 200 30m $\times$
HSIHybrid $\sim$4m 9x9 Multiple Variable Variable $\times$
HyperGlobal-450K $\sim$447k 64x64 EO-1 and Gaoen-5B 242 and 330 30m $\times$
SpectralEarth $\sim$538k 128x128 EnMAP 202 30m $\checkmark$
Table 1: Summary of Hyperspectral Remote Sensing Datasets: SpectralEarth contains close to nine gigapixels of 202 EnMAP bands each at 30m spatial resolution. This equates to an area of a little over six million square kilometers. The rows above the dashed line concern manually annotated datasets, while the part below lists unlabeled datasets for SSL.
Traditionally, hyperspectral benchmark datasets have been limited in geographical coverage, often confined to a single or a limited number of scenes [60, 3, 49, 30]. This limitation primarily roots in the expenses associated with airborne hyperspectral sensor surveys on the one hand, and the collection of ground truth annotations and in-situ measurements on the other hand. While the collection of quality labels remains a challenge today, the increased accessibility to more cost-effective sensors, and the launch of hyperspectral satellite missions enabled coverage of larger geospatial regions by HSI. Developments in SSL and foundation models have also driven interest in creating large-scale unlabeled HSI datasets. For instance, the HySpecNet-11k [19] collected about 11,500 unlabeled hyperspectral patches from the EnMAP satellite. HySpecNet-11k is designed to benchmark (unsupervised) image compression techniques. Similarly, MSST [55] compiled a dataset of 19,792 EnMAP patches for SSL pre-training. While HySpecNet-11k and MSST have contributed to increase the scale of existing HSI datasets, they remain limited in size to train foundation models. The recently introduced HyperGlobal-450K dataset [63] encompasses 450K HSI patches derived from 486 Hyperion EO-1 and 215 Gaofen-5B scenes. While this dataset significantly increases the scale of available HSI data, it presents the following limitations: First, its geographical distribution is less diverse, with a substantial part of the data concentrated in a few provinces of China. Moreover, it relies on imagery from the decommissioned EO-1 satellite. In contrast, SpectralEarth offers more extensive global coverage using data from the currently operational EnMAP satellite continuously collecting data. In addition, SpectralEarth is approximately five times larger in pixel count and it includes a temporal dimension increasing its value for pre-training of hyperspectral foundation models. A comparison of existing labeled and unlabeled HSI datasets provides Table 1.
2.2 Self-Supervised Learning for Remote Sensing
@@
In remote sensing, SSL techniques have been applied and adapted to multispectral and high-resolution RGB imagery. Manas et al. proposed SeCo [46], which compiled a 200K-location dataset comprising Sentinel 2 images with multiple temporal views (one per season) and introduced seasonal contrast augmentation. Their results supported the benefit of incorporating seasonal invariance in the training objective of contrastive methods. Building on this approach, Wang et al. proposed SSL4EO-S12 [65], which provided co-located Sentinel 2 and Sentinel 1 images and pre-trained models for both sensors. Subsequently, Stewart et al. extended this approach to Landsat imagery with SSL4EO-L [56], further expanding the available pre-training data, downstream datasets, and pre-trained models. Several studies have adapted MAE for remote sensing imagery. SatMAE [16] tokenized inputs across spectral and temporal dimensions, while Scale-MAE [54] introduced ground-sampled distance positional encoding for multi-scale imagery. GFM [47] assembled GeoPile, a pre-training dataset comprising high-resolution RGB images, by merging a variety of existing datasets. GeoPile was used to train an MAE model with an additional distillation loss guided by an ImageNet-22K pre-trained teacher. SpectralGPT [31] and S2MAE [42] proposed a hierarchical MAE pre-training approach based on a spatial-spectral vision transformer for Sentinel 2 imagery. CROMA [20] combined contrastive learning and MIM for SAR-Optical self-supervised learning. msGFM [23] proposed a multi-sensor MAE model with a cross-sensor reconstruction loss for paired images on RGB, Sentinel 1, Sentinel 2, and DSM data.
@@
In the hyperspectral domain, most SSL works focus on pixel-level classification (with or without spatial context) where the pre-training scene is the same as the test image [43, 32, 44, 9, 10, 33, 59]. While this approach is effective for classical HSI datasets, the learned representations cannot generalize across scenes. To address this issue, HSIMAE [66] proposed HSIHybrid, a compilation of existing HSI datasets, and pre-trained an MAE-like model on it. While this approach effectively increases data diversity, it poses challenges when combining sensors with different characteristics, requiring pre-processing steps such as PCA to unify the number of spectral bands. Additionally, the scale of HSIHybrid remains limited for training HSI foundation models. Recent studies have explored collecting unlabeled HSI data for pre-training. For example, MSST [55] used MAE on EnMAP imagery, introducing decoupled spectral and spatial attention blocks to reduce the computational cost associated with spatial-spectral tokenization. HyperSIGMA [63] proposed an MAE-inspired approach with independent spectral and spatial transformers and evaluated their model on several hyperspectral downstream tasks. However, the data used for pre-training these models is less diverse than SpectralEarth. Additionally, these studies focus exclusively on ViTs and MAE, excluding ConvNets and other pre-training methodologies. In contrast, our work explores a more diverse set of architectures and pre-training strategies and provides a library of pretrained models for HSI applications.
3 SpectralEarth & Benchmarks
@@
This section introduces the SpectralEarth dataset and related downstream datasets for benchmarking. Additional details are provided in Appendix.
(a) Classes: Arable land, Coniferous forest, Moors, heathland, sclerophyllous vegetation.
(b) Classes: Urban fabric, Industrial units, Arable land, Natural grassland, sparsely vegetated areas.
(c) Classes: Urban fabric, Arable land, Permanent crops, Complex cultivation patterns, Inland waters.
(d) Classes: Urban fabric, Arable land, Complex cultivation patterns, Coniferous forest.
Figure 4: Sample pseudo-RGB images of the curated EnMAP-CORINE multi-label classification benchmark.
(a) EnMAP-CDL
(b) EnMAP-NLCD
Figure 5: Sample image-mask (left-right) pairs of the curated EnMAP-CDL (a) and EnMAP-NLCD (b) semantic segmentation benchmarks.
3.1 SpectralEarth
Data Acquisition.

### v2

@@ Abstract
Foundation models have triggered a paradigm shift in computer vision and are increasingly being adopted in remote sensing, particularly for multispectral imagery. Yet, their potential in hyperspectral imaging (HSI) remains untapped due to the absence of comprehensive and globally representative hyperspectral datasets. To close this gap, we introduce SpectralEarth, a large-scale multi-temporal dataset designed to pretrain hyperspectral foundation models leveraging data from the Environmental Mapping and Analysis Program (EnMAP). SpectralEarth comprises 538,974 image patches covering 415,153 unique locations from 11,636 globally distributed EnMAP scenes spanning two years of archive. Additionally, 17.5% of these locations include multiple timestamps, enabling multi-temporal HSI analysis. Utilizing state-of-the-art self-supervised learning (SSL) algorithms, we pretrain a series of foundation models on SpectralEarth, integrating a spectral adapter into classical vision backbones to accommodate the unique characteristics of HSI. In tandem, we construct nine downstream datasets for land-cover, crop-type mapping, and tree-species classification, providing benchmarks for model evaluation. Experimental results support the versatility of our models and their generalizability across different tasks and sensors. We also highlight computational efficiency during model fine-tuning. The dataset, pretrained models, and code are publicly available. [^1] Dataset: https://geoservice.dlr.de/web/datasets/enmap_spectralearth, Code: https://github.com/AABNassim/spectral_earth
footnotetext: This is the accepted manuscript version of the article published in IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing. The final version is available at: https://doi.org/10.1109/JSTARS.2025.3581451
I Introduction
@@
Hyperspectral imaging (HSI) from space provides valuable information about the material composition of the Earth’s surface and atmosphere. With hundreds of narrow bands, each a few nanometers in width, hyperspectral images record detailed electromagnetic information across a wide range of wavelengths, from long-wave ultraviolet to short-wave infrared [1]. This rich spectral information provides opportunities for numerous environmental applications such as soil and mineral mapping, pollution tracking, agricultural assessment, and forest monitoring [2, 3, 4, 5, 6].
@@
Figure 1: Visualization of scale for various hyperspectral datasets published within the past three decades illustrating the volume of the SpectralEarth dataset (area of circles).
@@
In recent years, the availability of hyperspectral data has been significantly expanded by the launch of new satellite missions. Notable examples include Germany’s Environmental Mapping and Analysis Program (EnMAP) [7] with precursor DLR Earth Sensing Imaging Spectrometer mission (DESIS) [8] mounted to the International Space Station (ISS), Italy’s Hyperspectral Precursor and Application Mission (PRISMA) [9], and the European Space Agency’s upcoming Copernicus Hyperspectral Imaging Mission for the Environment (CHIME) [10]. These developments open new avenues to employ deep learning and Self-Supervised Learning (SSL) for large-scale HSI analysis.
@@
Figure 2: Mosaic of pseudo-RGB representatives from the SpectralEarth dataset to showcase various landscapes, including urban areas, agricultural land, deserts, forests, and water bodies. Each image is presented with a false-color composite using the first three principal components.
@@
Foundation models pretrained with SSL demonstrated remarkable generalization capabilities with minimal fine-tuning in computer vision [11]. Subsequently, there has been a surge in developing geospatial foundation models, particularly for multispectral and high-resolution RGB imagery [12]. This line of research benefits from the abundance of publicly available satellite data from missions like Copernicus Sentinel-2, which provides petabytes of archive data for model training. In contrast, progress in foundation models for the hyperspectral domain has been hindered by the lack of large-scale HSI datasets for pre-training. Historically, access to hyperspectral imagery has been costly, with most benchmarks restricted to small-scale aerial imagery datasets [13]. Although recent efforts [14, 15, 16] have begun to address this limitation, they still fall short in terms of data volume and geographic diversity to effectively pre-train general-purpose hyperspectral models.
@@
To bridge this gap, we introduce SpectralEarth, a large-scale dataset derived from the EnMAP satellite mission, featuring over 538,974 hyperspectral image patches from 415,153 unique locations with global spatial distribution and cloud coverage below 10%. SpectralEarth, as illustrated in Figure 1, represents an important leap in scale, being significantly larger than existing HSI datasets, and spans a wide variety of landscapes to reflect the full diversity of spectral signatures. Compiled from 11,636 EnMAP scenes, the dataset volume exceeds 3TB in hyperspectral images. Notably, about 17.5% of its geospatial locations include time series data to enable multi-temporal HSI analysis. To harness the rich information in hyperspectral data, we modify conventional vision backbones such as ResNet [17] and ViT [18] with a spectral adapter module. This enables such backbones to capture the unique spectral characteristics of HSI. Using three popular SSL algorithms, MoCo-V2 [19, 20], DINO [21] and MAE [22], we train these backbones on SpectralEarth, to provide a plurality of pretrained models for hyperspectral image analysis. To benchmark the efficacy of our models, we introduce nine downstream datasets constructed by pairing EnMAP, DESIS, and Hyperion EO-1 imagery with land cover, crop type, and tree species labels. Our experiments demonstrate the potential of large-scale pre-training to efficiently generalize across various hyperspectral imaging contexts, including adapting to data from alternative sensors with different spectral characteristics. Our contributions are summarized as follows:
@@
  • Assembly of SpectralEarth, a large-scale dataset comprising around 3.3TB of nearly cloud-free EnMAP hyperspectral images. SpectralEarth features a global geospatial distribution encompassing over 538,974 image patches from 415,153 unique geo-locations, 73,307 including multiple timestamps.
@@
  • Construction of nine downstream datasets designed to benchmark classification and semantic segmentation tasks for HSI.
@@
  • An empirical evaluation of self-supervised learning in hyperspectral imaging, by pre-training various models on SpectralEarth and evaluating their generalizability across downstream tasks and sensors. In addition, we demonstrate the models’ computational benefit through faster convergence in fine-tuning.
II Related Work
@@ II-A Hyperspectral Datasets
Indian Pines [23] 1 $\times145$ AVIRIS 224 20m $\times$
@@
[…]
Pavia Center [24] 1 $\times1096$ ROSIS 102 1.3m $\times$
[…]
SDFC [26] 1 $1×\times20$1 HAHS 63 0.5m $\times$
[…]
Hyperview [29] $\sim$3k avg. $\sim$$\times60$ HySpex 150 2m $\times$
[…]
HySpecNet-11k [14] $\sim$11k $\times128$ EnMAP 202 30m $\times$
MSST [15] $\sim$20k $\times64$ EnMAP 200 30m $\times$
HSIHybrid [31] $\sim$4m $\times9$ Multiple Variable Variable $\times$
HyperGlobal-450K [16] $\sim$447k $\times64$ EO-1 and Gaofen-5B 242 and 330 30m $\times$
SpectralEarth $\sim$538k $\times128$ EnMAP 202 30m $\checkmark$
TABLE I: Summary of Hyperspectral Remote Sensing Datasets: SpectralEarth contains nearly nine gigapixels of 202 EnMAP bands each at 30m spatial resolution. This equates to an area of a little over six million square kilometers. The rows above the dashed line concern manually annotated datasets, while the part below lists unlabeled datasets for SSL.
Traditionally, hyperspectral benchmark datasets have been limited in geographical coverage, often confined to a single or a limited number of scenes [32, 13, 29, 33]. This limitation is primarily rooted in the expenses associated with airborne hyperspectral sensor surveys on the one hand and the collection of ground truth annotations and in-situ measurements on the other hand. While collecting quality labels remains a challenge today, the increased accessibility to more cost-effective sensors and the launch of hyperspectral satellite missions enabled coverage of larger regions with HSI. Developments in SSL and foundation models have also driven interest in creating large-scale unlabeled HSI datasets. For instance, the HySpecNet-11k [14] collected about 11,500 unlabeled hyperspectral patches from the EnMAP satellite. HySpecNet-11k is designed to benchmark (unsupervised) image compression techniques. Similarly, MSST [15] compiled a dataset of 19,792 EnMAP patches for SSL pre-training. While HySpecNet-11k and MSST have contributed to increasing the scale of existing HSI datasets, they remain limited in size to train foundation models. The recently introduced HyperGlobal-450K dataset [16] encompasses more than 450,000 HSI patches derived from 486 Hyperion EO-1 and 215 Gaofen-5B scenes. While this dataset significantly increases the scale of publicly available HSI data, it presents certain limitations. First, its geographical distribution is less diverse, with a substantial part of the data concentrated in a few provinces of China. Moreover, it relies on imagery from the decommissioned EO-1 satellite. In contrast, SpectralEarth offers more extensive global coverage using data from the currently operational EnMAP satellite, which is continuously collecting data. In addition, SpectralEarth is approximately five times larger in pixel count and includes a temporal dimension, increasing its value for pre-training of hyperspectral foundation models. A comparison of existing labeled and unlabeled HSI datasets is provided in Table I.
II-B Self-Supervised Learning for Remote Sensing
@@
In remote sensing, SSL techniques have been applied and adapted to multispectral and high-resolution RGB imagery. SeCo [47] compiled a 200K-location dataset comprising Sentinel 2 images with multiple temporal views (one per season) and introduced seasonal contrast augmentation. The results supported the benefit of incorporating seasonal invariance in the training objective of contrastive methods. Building on this approach, SSL4EO-S12 [48] provided co-located Sentinel 2 and Sentinel 1 images and pre-trained models for both sensors. Subsequently, SSL4EO-L [49] extended this approach to Landsat imagery, further expanding the available pre-training data, downstream datasets, and pre-trained models. Several studies have adapted MAE for remote sensing imagery. SatMAE [50] tokenized inputs across spectral and temporal dimensions, while Scale-MAE [51] introduced ground-sampled distance positional encoding for multi-scale imagery. GFM [52] assembled GeoPile, a pre-training dataset comprising high-resolution RGB images, by merging various existing datasets. GeoPile was used to train an MAE model with an additional distillation loss guided by an ImageNet-22K pre-trained teacher. SpectralGPT [53] and S2MAE [54] proposed a hierarchical MAE pre-training approach based on a spatial-spectral vision transformer for Sentinel 2 imagery. Prithvi [55] and Prithvi-EO-2.0 [56] built large-scale pretraining datasets from Harmonized Landsat-Sentinel-2 time-series and spatial/spectral/temporal ViT backbones using MAE. OmniSat [57] proposed a multi-sensor framework that combines masked autoencoding with contrastive learning, using high-resolution aerial imagery and multi-modal time series from Sentinel-1 and Sentinel-2. CROMA [58] combined contrastive learning and MIM for SAR-Optical self-supervised learning. msGFM [59] proposed a multi-sensor MAE model with a cross-sensor reconstruction loss for paired images on RGB, Sentinel 1, Sentinel 2, and DSM data. SenPa-MAE [60] introduced a sensor-aware MAE pretraining strategy by incorporating metadata such as ground sampling distance and spectral response functions to allow sensor generalization. The model was trained with MAE on Sentinel-2, Planet-SuperDove, and Landsat-8/9. AnySat [61] leveraged I-JEPA [62] to learn joint representations across several sensors. The model was trained on a combination of multiple datasets, including high-resolution imagery, Sentinel 1 and 2, MODIS, Landsat-6/7/8, and ALOS-2 time series.
@@
In the hyperspectral domain, most SSL works focus on pixel-level classification (with or without spatial context) where the pre-training scene is the same as the test image [63, 64, 65, 66, 67, 68, 69]. While this approach is effective for classical HSI datasets, the learned representations cannot generalize across scenes. To address this issue, HSIMAE [31] proposed HSIHybrid, a compilation of existing HSI datasets, and pre-trained an MAE-like model. This approach effectively increases data diversity but poses challenges when combining sensors with different characteristics, requiring pre-processing steps such as PCA to unify the number of spectral bands. Additionally, the scale of HSIHybrid remains limited when training HSI foundation models. Recent studies have explored collecting unlabeled HSI data for pre-training. For example, MSST [15] used MAE on EnMAP imagery, introducing decoupled spectral and spatial attention blocks to reduce the computational cost associated with spatial-spectral tokenization. DOFA [70] proposed a general-purpose architecture that dynamically generates patch embedding weights from sensor wavelength metadata. The model was trained using MAE combined with a distillation loss on RGB, multispectral, SAR, and EnMAP hyperspectral imagery. HyperSIGMA [16] proposed an MAE-inspired approach with independent spectral and spatial transformers and evaluated their model on several hyperspectral downstream tasks. However, the data used for pre-training these models is less diverse than SpectralEarth. Additionally, these studies focus exclusively on ViTs and MAE, excluding ConvNets and other pre-training methodologies. In contrast, our work explores a more diverse set of architectures and pre-training strategies and provides a library of pretrained models for HSI applications.
III SpectralEarth & Benchmarks
@@
This section introduces the SpectralEarth dataset and related downstream datasets for benchmarking.
(a) CORINE
(b) CDL
(c) NLCD
(d) TreeMap
(e) BD-Foret
(f) EuroCrops
(g) BNETD
Figure 4: Sample pseudo-RGB images from the proposed EnMAP benchmarks.
III-A SpectralEarth
Data Acquisition